OpenStudy (anonymous):

HELP PLEASE p x 2300 289 2000 302 1700 312 1500 325 1300 334 1200 342 1000 359 p=a+b lnx whats a and b

OpenStudy (anonymous):

If you use p=a+blnx, in the first row use: 2300=a+bln289 now solve for a once you have that a use the next row: 2000=a+bln302 plug the a you solve for up top in this equation, then you can solve for b, once you know b plug that back into your first equation to get a, does that make sense?

OpenStudy (anonymous):

.... no sorry

OpenStudy (anonymous):

don't let them fool you, lnx is just a number for this exercise don't look at it as a variable

OpenStudy (anonymous):

im so lost...

OpenStudy (anonymous):

So we have two equations: 1 -> 2300=a+bln289 2 -> 2000=a+bln302 If we have 2 equations and 2 unknowns we can solve one then substitute it into the other. Look at the equations like this: (this has the ln solved for) 1 -> 2300=a+b (5.666) 2 -> 2000=a+b (5.71) Get equation 1 to be a = .... Then plug that into equation 2 so you will only have b terms, then you can solve for a number for b, and plug back into 1 so you will only have a to solve for.

OpenStudy (anonymous):

so a = 406?

OpenStudy (anonymous):

Is your equation p=(a+b) lnx or p=a+(b lnx)? for the initial problem we are looking at?

OpenStudy (anonymous):

the second 1 i think

OpenStudy (anonymous):

So then... 1 -> 2300=a+b (5.666) 2 -> 2000=a+b (5.71) Going from here, solving for a in 1 i get: a=2300-(b*5.666) Then i can plug this into the 2nd equation so: 2000= (2300-(b*5.666))+(b*5.71) If you solve for b here I get b=~-6818 plug this back into: a=2300-(b*5.666) a=2300-(-6818*5.666) a=~40,930 Does that make sense? Honestly to me those numbers seem a little strange, but if the equation is setup the second way like you said that is what I got going through the math.

OpenStudy (anonymous):

thx :)